### v2

@@ 1.2 Related work
We have reviewed the works most closely related to our memory capacity results, but there are many relevant studies exploring the optimization of one-layer transformers as well, such as Tian et al (2023), Chen and Li (2024), Tarzanagh et al (2023), Zhang et al (2024), and Deora et al (2024). However, most do not address the next-token prediction scenario we study here. The exceptions, such as Makkuva et al (2024), Li et al (2024), Tian et al (2023), Tian et al (2024), and Ildiz et al (2024), rely on restrictive assumptions about input data and transformer dimensions, such as embedding size, to achieve minimal empirical risk. Additionally, optimization and learnability in the next-token prediction setting, albeit limited to linear models, have been recently studied in Malach (2023) and Thrampoulidis (2024). Finally, the effectiveness of transformers for particular tasks has been explored in papers such as Sanford et al (2023) and Wang et al (2024).
Miscellaneous results.
@@
Our upper bound, Theorem 4.5, can be seen as a generalization of Theorem 1 of Sontag (1997) from input-output pairs in $\mathbb{R}^{d}\times\mathbb{R}$ to input-output pairs in $\cup_{t=1}^{\infty}[\omega]^{t}\times\Delta^{\omega-1}$. However, our proof is quite different because $\cup_{t=1}^{\infty}[\omega]^{t}$ is countable. As for Theorems 6.3 and 6.4, these are classical results when $a$ and $b$ are both generic, but to the best of our knowledge, they are novel results in the setting where $b\in\mathbb{R}^{n}$ has entries which are nonzero and distinct. However, closely related results have been shown before. Tamura (1991) proved a version of Theorem 6.4 when $\psi$ is the logistic function and restated the result in Tamura and Tateishi (1997). Then Huang and Babri (1998) proved it more generally when $\psi$ is bounded, nonlinear, and has a limit at either positive or negative infinity. Their result does not require $\psi$ to be real analytic at a point and not a polynomial there, but our result does not require $\psi$ to be bounded or to have limits. Thus, the two results are complementary. Interestingly, their proof sends the inputs of $\psi$ off to positive or negative infinity, while our proof only uses inputs in the interval of convergence of the point where $\psi$ is real analytic.
@@ 1.3 Organization
In Section 2 we present notations, conventions, and prerequisites. In Section 3 we carefully delineate the assumptions made to start talking about next-token prediction. Section 3 serves to ground the definition of next-token prediction capacity, but the reader that is already willing to accept the definition can safely skip to Section 4. In Section 4 we define next-token prediction capacity and prove the upper bound Theorem 4.5. In Section 5 we define the one-layer multi-head decoder-only transformer model. In Section 6 we prove the injectivity result, the rank results, and the next-token prediction lower bound. In Section 7 we present a numerical experiment demonstrating that not only is the memory capacity on the same order as the number of parameters divided by the vocabulary size, but the transformer can already be trained to the entropy lower bound in this regime. In Section 8 we introduce the token-averaged FNN model and prove a next-token prediction capacity lower bound for it. In Section 9, we conclude.
@@ 2 Preliminaries
In the realm of real analytic functions, we use Taylor’s theorem as as well as the fact that the zero set of a nontrivial real analytic function is measure zero and closed (Gunning and Rossi, 1965, Corollary 10). We say a property holds “generically” if the set of exceptions lies on the zero set of a nontrivial real analytic function. Note that $\phi$ is real analytic since the numerator and denominator are real analytic with the denominator being nowhere zero. Moreover, the entire self-attention sub-layer of the transformer model, which will be described in detail below, is real analytic since it is a composition of real analytic functions. Furthermore, if $\psi$ is real analytic at a point, then the entire transformer parameter mapping is real analytic at a point. Similarly, if $\psi$ is continuously differentiable, then the entire transformer parameter mapping is continuously differentiable, in which case we can apply Sard’s theorem (Sard, 1942), which states that the set of critical values has measure zero. A critical point is a point where the differential is not surjective and a critical value is a value with level set (preimage/fiber) containing at least one critical point.
@@
We also use the model-theoretic structure $(\mathbb{R}_{\mathrm{an}},\exp)$ (van den Dries et al, 1994; van den Dries and Miller, 1994; van den Dries, 1998). Polynomials, restricted analytic functions, and exp are the basic $(\mathbb{R}_{\mathrm{an}},\exp)$-definable functions, but compositions, quotients (if the denominator is nowhere zero), inverses (of injective functions), and limits of $(\mathbb{R}_{\mathrm{an}},\exp)$-definable functions are $(\mathbb{R}_{\mathrm{an}},\exp)$-definable as well (van den Dries, 1998, Lma. 2.3). From this, it can be seen that $\phi$, tanh, arctan, erf, and GELU are $(\mathbb{R}_{\mathrm{an}},\exp)$-definable. Furthermore, if $\psi$ is $(\mathbb{R}_{\mathrm{an}},\exp)$-definable, then the entire transformer parameter mapping is $(\mathbb{R}_{\mathrm{an}},\exp)$-definable, and so its image either has nonempty interior or is nowhere dense (Sontag, 1996, Corollary B.4).
@@
In the realm of linear algebra, we use the notion of Kruskal rank (Kruskal, 1977), as well as the Cauchy-Binet formula (Gantmacher, 1960, Sec. I.2.4) and the Leibniz determinant formula (Axler, 2015, Def. 10.33). In the realm of number theory, we use that $\exp(1)$ is transcendental (Hermite, 1873). Finally, we also use basic concepts from probability theory to which we refer the reader to Billingsley (1995) as a reference.
@@ 3 Probabilistic language space
Axiom 5 of the first definition says that $p(\cdot\mid x)$ is a conditional probability mass function whenever it is defined. Axiom 4 says that $p(\cdot\mid x)$ is defined for all $x$ except those which have zero probability by the chain rule. Definition 3.2 essentially applies the chain rule to the conditional probability mass functions from Definition 3.1. Note that for all $t\geq 2$, $\sum_{x\in\mathcal{X}^{t}}p(x)=\sum_{y\in\mathcal{X}^{t-1}}\sum_{z\in\mathcal{X}}p(y,z)=\sum_{y\in\mathcal{X}^{t-1}}p(y)$, so, by induction, $\sum_{x\in\mathcal{X}^{t}}p(x)=1$ for all $t\in\mathbb{N}$. In other words, $p$ is a probability mass function on $\mathcal{X}^{t}$ for every $t\in\mathbb{N}$, but also is compatible with being decomposed into the conditional probability mass functions of Definition 3.1.
The third definition requires greater explanation. Subsets of $\mathcal{X}^{\mathbb{N}}$ of the form $\mathcal{C}(A)$ for some $A\subset\mathcal{X}^{t}$ are called cylinders of rank $t$. Subsets of $\mathcal{X}^{\mathbb{N}}$ of the form $\mathcal{C}(\{x\})$ for some $x\in\cup_{t=1}^{\infty}\mathcal{X}^{t}$ are called thin cylinders. Let $\mathscr{C}_{0}$ be the class of cylinders. Then $\mathscr{C}_{0}$ is an algebra, as shown after Eq. (2.17) in Billingsley (1995). The $\sigma$-algebra generated by it, $\mathscr{C}$, is called the cylindrical $\sigma$-algebra of $\mathcal{X}^{\mathbb{N}}$. Axioms 3 and 4 of Definition 3.3 ensure that $P$ can be reduced to a $p$ as in Definition 3.2.
@@
Define $\Phi_{12}:\mathbb{L}_{1}\to\mathbb{L}_{2}$ in the following way. Let $p:\mathcal{X}\times A\to[0,1]$ be in $\mathbb{L}_{1}$. Extend $p$ to all of $\mathcal{X}\times\cup_{t=0}^{\infty}\mathcal{X}^{t}$ by setting $p(y\mid x)=0$ for all $(y,x)\in\mathcal{X}\times A^{\mathsf{c}}$. Then, for all $x\in\cup_{t=1}^{\infty}\mathcal{X}^{t}$, define $q(x)=\prod_{t=1}^{|x|}p(x_{t}\mid x_{1},\ldots,x_{t-1})$. Then $(\mathcal{X},q)$ satisfies the first three axioms of Definition 3.2 by construction. To see that it satisfies the fourth axiom as well, let $x\in\cup_{t=1}^{\infty}\mathcal{X}^{t}$ and observe
@@
We will show that $\Phi_{12}$ is a bijection by constructing its inverse. Define $\Phi_{21}:\mathbb{L}_{2}\to\mathbb{L}_{1}$ in the following way. Let $q\in\mathbb{L}_{2}$. Extend $q$ to all of $\cup_{t=0}^{\infty}\mathcal{X}^{t}$ by setting $q(\leavevmode\nobreak\ )=1$. Define $A=\mathcal{X}^{0}\cup\{x\in\cup_{t=1}^{\infty}\mathcal{X}^{t}\mid q(x)\neq 0\}$ and $p:\mathcal{X}\times A\to[0,1]:(y,x)\mapsto q(x,y)/q(x)$. Then $p$ is well-defined and satisfies the first four axioms of Definition 3.1 by construction. To see that it satisfies the fifth axiom, we will use induction. First, $\sum_{y\in\mathcal{X}}p(y\mid\leavevmode\nobreak\ )=\sum_{y\in\mathcal{X}}q(y)=1$. Now let $t\in\mathbb{N}$ and suppose $\sum_{y\in\mathcal{X}}p(y\mid x)=1$ for all $x\in A$ such that $|x|<t$. Then let $x\in A$ such that $|x|=t$ and observe
@@
Thus, by induction, $P_{0}$ is well-defined. Next, we will show that $P_{0}$ is a probability measure on $\mathscr{C}_{0}$. We already have $P_{0}(\emptyset)=0$. To get $P_{0}(\Omega)=1$ we can use $\Omega=\mathcal{C}(\mathcal{X})$ and $\sum_{x\in\mathcal{X}}p(x)=1$. So the only thing left to show is that $P_{0}$ is countably additive. But, by Theorem 2.3 of (Billingsley, 1995), the class of cylinders has the special property that finitely additive probability measures on it are countably additive as well. Thus, we only need to show that $P_{0}$ is finitely additive. Let $\mathcal{C}(A)$ and $\mathcal{C}(B)$ be disjoint. Since every cylinder of rank $t_{0}$ is also a cylinder of rank $t$ for all $t\geq t_{0}$, we can choose $A,B$ to be subsets from the same $\mathcal{X}^{t}$. Thus, $A$ and $B$ are disjoint and $\mathcal{C}(A)\cup\mathcal{C}(B)=\mathcal{C}(A\cup B)$. So,
$P_{0}(\mathcal{C}(A\cup B))=\sum_{x\in A\cup B}p(x)=\sum_{x\in A}p(x)+\sum_{x\in B}p(x)=P_{0}(\mathcal{C}(A))+P_{0}(\mathcal{C}(B)),$
@@
proving that $P_{0}$ is finitely additive, hence countably additive, hence a probability measure on $\mathscr{C}_{0}$. So, by the Extension Theorem (Fréchet, 1924), $P_{0}$ has a unique extension to a probability measure $P$ on $\mathscr{C}$. Define $\Phi_{23}(p)=P$.
@@
It is obvious that $(\Phi_{32}\circ\Phi_{23})(p)=p$ for all $p\in\mathbb{L}_{2}$. To prove that $(\Phi_{23}\circ\Phi_{32})(P)=P$ for all $P\in\mathbb{L}_{3}$, on the other hand, is quite subtle. Let $P_{1}\in\mathbb{L}_{3}$. Define $p=\Phi_{32}(P_{1})$ and $P_{2}=\Phi_{23}(p)$. While the uniqueness part of the Extension Theorem allowed us to define $\Phi_{23}$, it does not guarantee that $P_{1}=P_{2}$. To prove the latter, we will use Theorem 3.3 of (Billingsley, 1995), a consequence of Dynkin’s $\pi$-$\lambda$ theorem, which can be used to prove the uniqueness part of the Extension Theorem as well. Let $\mathscr{G}$ be the class of the empty set and thin cylinders. It is easy to check that $\mathscr{G}$ is closed under binary intersections and is, therefore, a $\pi$-system. Also, every cylinder can be decomposed into a finite union of thin cylinders. Thus, the algebra generated by $\mathscr{G}$ is $\mathscr{C}_{0}$, hence the $\sigma$-algebra generated by $\mathscr{G}$ is $\mathscr{C}$. So, by Theorem 3.3 of (Billingsley, 1995), any two probability measures on $\mathscr{C}$ that agree on $\mathscr{G}$ also agree on $\mathscr{C}$. Let $x\in\cup_{t=1}^{\infty}\mathcal{X}^{t}$. Then $P_{2}(\mathcal{C}(\{x\}))=p(x)=P_{1}(\mathcal{C}(\{x\}))$. So, $P_{1}$ and $P_{2}$ agree on $\mathscr{G}$, and so they are equal. Thus, $\Phi_{23}$ is the inverse of $\Phi_{32}$. ∎
@@
Note that the mapping $\Phi_{23}$ is standard. It follows the steps in Billingsley (1995) almost exactly until the definition of $P_{0}$ where Billingsley (1995) instead defines the product measure, which is the independent version of $P_{0}$. But, in the dependent setting, our $P_{0}$ is also standard, see e.g. Eq. (8) of Du et al (2023), as it is the only possible definition that satisfies finite additivity since $\mathcal{C}(A)=\cup_{x\in A}\mathcal{C}(\{x\})$. The final step of extending $P_{0}$ to $P$ also follows as in Billingsley (1995). But the subtle point that we have made in Proposition 3.5 is that for every $p\in\mathbb{L}_{2}$ there is a unique $P\in\mathbb{L}_{3}$ such that $P(\mathcal{C}(\{x\}))=p(x)$ for all $x\in\cup_{t=1}^{\infty}\mathcal{X}^{t}$. While the Extension Theorem guarantees that the extension from $\mathscr{C}_{0}$ to $\mathscr{C}$ is unique, which implies that $\Phi_{23}$ is well-defined, it does not prove that the extension from $\mathscr{G}$ to $\mathscr{C}$ is unique. So, while previous works showed that $\Phi_{23}$ is well-defined, they did not seem to notice that it is a bijection.
@@
Propositions 3.4 and 3.5 have important philosophical implications. Specifically, assuming that the next word in a sequence of words is always governed by a probability distribution is equivalent to assuming that infinite sequences of words are governed by a distribution. Moreover, the underlying functions in the two cases are in one-to-one correspondence with one another. But, it is conceptually easier to believe that there is an underlying probabilistic language space of the first kind because we naturally think of a sequence of words along with its extra-mathematical context, such as body language and flowers along the road. On the other hand, an underlying probabilistic language space of the third kind implies that language is a stochastic process. To see this, note that every probability measure is the pushforward measure of some random variable: just let $X$ be the identity mapping from $(\Omega,\mathscr{F})$ to itself, then $PX^{-1}=P$. In particular, for every $P\in\mathbb{L}_{3}$, there is a probability space $(\Omega,\mathscr{F},\mu)$ and a $(\mathcal{X}^{\mathbb{N}},\mathscr{C})$-valued random variable $X$ on $(\Omega,\mathscr{F},\mu)$ such that $P=\mu X^{-1}$. Each $X_{t}$ is a $(\mathcal{X},2^{\mathcal{X}})$-valued random variable, so $X$ is a stochastic process. Furthermore, its finite-dimensional distributions $\mu(X_{1},\ldots,X_{t})^{-1}$ are given by the probability mass function $\Phi_{32}(P)$ for each $t\in\mathbb{N}$.

### v3

@@ 1.2.3 Optimization of transformers
We have reviewed the works most closely related to our memory capacity results, but there are many relevant studies exploring the optimization of one-layer transformers as well, such as [27, 28, 29, 30, 31]. However, most do not address the next-token prediction scenario we study here. The exceptions, such as [32, 33, 27, 34, 35], rely on restrictive assumptions about input data and transformer dimensions, such as embedding size, to achieve minimal empirical risk. Additionally, optimization and learnability in the next-token prediction setting, albeit limited to linear models, have been recently studied in [36, 37]. Finally, the effectiveness of transformers for particular tasks has been explored in papers such as [38, 39].
1.2.4 Miscellaneous results
@@
Our upper bound, Theorem 13, can be seen as a generalization of Theorem 1 of Sontag [40] from input-output pairs in $\mathbb{R}^{d}\times\mathbb{R}$ to input-output pairs in $\cup_{t=1}^{\infty}[\zeta]^{t}\times\Delta^{\zeta-1}$. However, our proof is quite different because $\cup_{t=1}^{\infty}[\zeta]^{t}$ is countable. As for Theorems 19 and 20, these are classical results when $a$ and $b$ are both generic, but to the best of our knowledge, they are novel results in the setting where $b\in\mathbb{R}^{n}$ has entries which are nonzero and distinct. However, closely related results have been shown before. Tamura [41] proved a version of Theorem 20 when $\psi$ is the logistic function and restated the result in [42]. Then Huang and Babri [43] proved it more generally when $\psi$ is bounded, nonlinear, and has a limit at either positive or negative infinity. Their result does not require $\psi$ to be real analytic at a point and not a polynomial there, but our result does not require $\psi$ to be bounded or to have limits. Thus, the two results are complementary. Interestingly, their proof sends the inputs of $\psi$ off to positive or negative infinity, while our proof only uses inputs in the interval of convergence of the point where $\psi$ is real analytic.
@@ 1.3 Organization
In Section 2 we present notations, conventions, and prerequisites. In Section 3 we carefully delineate the assumptions made to start talking about next-token prediction. Section 3 serves to ground the definition of next-token prediction capacity, but the reader that is already willing to accept the definition can safely skip to Section 4. In Section 4 we define next-token prediction capacity and prove the upper bound Theorem 13. In Section 5 we define the one-layer multi-head decoder-only transformer model. In Section 6 we prove the injectivity result, the rank results, and the next-token prediction lower bound. In Section 7 we present a numerical experiment demonstrating that not only is the memory capacity on the same order as the number of parameters divided by the vocabulary size, but the transformer can already be trained to the entropy lower bound in this regime. In Section 8 we introduce the token-averaged FNN model and prove a next-token prediction capacity lower bound for it. In Section 9, we conclude.
@@ 2 Preliminaries
In the realm of real analytic functions, we use Taylor’s theorem as as well as the fact that the zero set of a nontrivial real analytic function is measure zero and closed [14, Corollary 10]. We say a property holds “generically” if the set of exceptions lies on the zero set of a nontrivial real analytic function. Note that $\phi$ is real analytic since the numerator and denominator are real analytic with the denominator being nowhere zero. Moreover, the entire self-attention sub-layer of the transformer model, which will be described in detail below, is real analytic since it is a composition of real analytic functions. Furthermore, if $\psi$ is real analytic at a point, then the entire transformer parameter mapping is real analytic at a point. Similarly, if $\psi$ is continuously differentiable, then the entire transformer parameter mapping is continuously differentiable, in which case we can apply Sard’s theorem [44].
Lemma 1 (Sard’s theorem).
Let $k\geq 1$. Let $M$ and $N$ be $C^{k}$ manifolds of dimension $m$ and $n$ respectively. Let $F:M\to N$ be $C^{k}$. If $m\leq n$ or $m\leq n+k-1$, then the set of critical values of $F$ has measure zero in $N$.
A critical point is a point where the differential is not surjective and a critical value is a value with level set (preimage/fiber) containing at least one critical point. For example, if $m<n$, then every point is a critical point, hence the image of $F$ has measure zero in $N$.
@@
We also use the model-theoretic structure $(\mathbb{R}_{\mathrm{an}},\exp)$ [45, 46, 47]. Polynomials, restricted analytic functions, and exp are the basic $(\mathbb{R}_{\mathrm{an}},\exp)$-definable functions, but compositions, quotients (if the denominator is nowhere zero), inverses (of injective functions), and limits of $(\mathbb{R}_{\mathrm{an}},\exp)$-definable functions are $(\mathbb{R}_{\mathrm{an}},\exp)$-definable as well [47, Lma. 2.3]. From this it can be seen that $\phi$, tanh, arctan, erf, and GELU are $(\mathbb{R}_{\mathrm{an}},\exp)$-definable. Furthermore, if $\psi$ is $(\mathbb{R}_{\mathrm{an}},\exp)$-definable, then the entire transformer parameter mapping is $(\mathbb{R}_{\mathrm{an}},\exp)$-definable, and so its image either has nonempty interior or is nowhere dense [48, Corollary B.4].
@@
In the realm of linear algebra, we use the notion of Kruskal rank [49], as well as the Cauchy-Binet formula [50, Sec. I.2.4] and the Leibniz determinant formula [51, Def. 10.33]. In the realm of number theory, we use that $\exp(1)$ is transcendental [52]. Finally, we also use basic concepts from probability theory to which we refer the reader to [11] as a reference.
For variables and functions, we will use the notation that is most appropriate to the section. However, there are some notations that are used consistently throughout the paper, so we will summarize those here. We use $\zeta$ to denote the number of tokens in the vocabulary. We use $\alpha,\beta,\gamma$ as example sequences from $\cup_{t=0}^{\infty}[\zeta]^{t}$. We use $N$ to denote the number of document sequences and $n$ to denote the number of distinct context sequences. We use $\theta$ to denote the concatenation of model parameters and $h_{\theta}$ to denote the model. We use $\phi$ to denote the softmax function and $\psi$ to denote the activation function. We use $Z$ for the embedding matrix, $d$ for the embedding dimension, $U$ for the position matrix, $m_{0}$ for the number of self-attention heads, $W$ and $V$ for self-attention and FNN matrices, and $m$ for the FNN width.
@@ 3 Probabilistic Language Space
Axiom 5 of the first definition says that $p(\cdot\mid x)$ is a conditional probability mass function whenever it is defined. Axiom 4 says that $p(\cdot\mid x)$ is defined for all $x$ except those which have zero probability by the chain rule. Definition 3 essentially applies the chain rule to the conditional probability mass functions from Definition 2. Note that for all $t\geq 2$, $\sum_{x\in\mathcal{X}^{t}}p(x)=\sum_{y\in\mathcal{X}^{t-1}}\sum_{z\in\mathcal{X}}p(y,z)=\sum_{y\in\mathcal{X}^{t-1}}p(y)$, so, by induction, $\sum_{x\in\mathcal{X}^{t}}p(x)=1$ for all $t\in\mathbb{N}$. In other words, $p$ is a probability mass function on $\mathcal{X}^{t}$ for every $t\in\mathbb{N}$, but also is compatible with being decomposed into the conditional probability mass functions of Definition 2.
The third definition requires greater explanation. Subsets of $\mathcal{X}^{\mathbb{N}}$ of the form $\mathcal{C}(A)$ for some $A\subset\mathcal{X}^{t}$ are called cylinders of rank $t$. Subsets of $\mathcal{X}^{\mathbb{N}}$ of the form $\mathcal{C}(\{x\})$ for some $x\in\cup_{t=1}^{\infty}\mathcal{X}^{t}$ are called thin cylinders. Let $\mathscr{C}_{0}$ be the class of cylinders. Then $\mathscr{C}_{0}$ is an algebra, as shown after Eq. (2.17) in [11]. The $\sigma$-algebra generated by it, $\mathscr{C}$, is called the cylindrical $\sigma$-algebra of $\mathcal{X}^{\mathbb{N}}$. Axioms 3 and 4 of Definition 4 ensure that $P$ can be reduced to a $p$ as in Definition 3.
@@
Define $\Phi_{12}:\mathbb{L}_{1}\to\mathbb{L}_{2}$ in the following way. Let $p:\mathcal{X}\times A\to[0,1]$ be in $\mathbb{L}_{1}$. Extend $p$ to all of $\mathcal{X}\times\cup_{t=0}^{\infty}\mathcal{X}^{t}$ by setting $p(y\mid x)=0$ for all $(y,x)\in\mathcal{X}\times A^{\mathsf{c}}$. Then, for all $x\in\cup_{t=1}^{\infty}\mathcal{X}^{t}$, define $q(x)=\prod_{t=1}^{|x|}p(x_{t}\mid x_{1},\ldots,x_{t-1})$. Then $(\mathcal{X},q)$ satisfies the first three axioms of Definition 3 by construction. To see that it satisfies the fourth axiom as well, let $x\in\cup_{t=1}^{\infty}\mathcal{X}^{t}$ and observe
@@
We will show that $\Phi_{12}$ is a bijection by constructing its inverse. Define $\Phi_{21}:\mathbb{L}_{2}\to\mathbb{L}_{1}$ in the following way. Let $q\in\mathbb{L}_{2}$. Extend $q$ to all of $\cup_{t=0}^{\infty}\mathcal{X}^{t}$ by setting $q(\leavevmode\nobreak\ )=1$. Define $A=\mathcal{X}^{0}\cup\{x\in\cup_{t=1}^{\infty}\mathcal{X}^{t}\mid q(x)\neq 0\}$ and $p:\mathcal{X}\times A\to[0,1]:(y,x)\mapsto q(x,y)/q(x)$. Then $p$ is well-defined and satisfies the first four axioms of Definition 2 by construction. To see that it satisfies the fifth axiom, we will use induction. First, $\sum_{y\in\mathcal{X}}p(y\mid\leavevmode\nobreak\ )=\sum_{y\in\mathcal{X}}q(y)=1$. Now let $t\in\mathbb{N}$ and suppose $\sum_{y\in\mathcal{X}}p(y\mid x)=1$ for all $x\in A$ such that $|x|<t$. Then let $x\in A$ such that $|x|=t$ and observe
@@
Thus, by induction, $P_{0}$ is well-defined. Next, we will show that $P_{0}$ is a probability measure on $\mathscr{C}_{0}$. We already have $P_{0}(\emptyset)=0$. To get $P_{0}(\Omega)=1$ we can use $\Omega=\mathcal{C}(\mathcal{X})$ and $\sum_{x\in\mathcal{X}}p(x)=1$. So the only thing left to show is that $P_{0}$ is countably additive. But, by Theorem 2.3 of [11], the class of cylinders has the special property that finitely additive probability measures on it are countably additive as well. Thus, we only need to show that $P_{0}$ is finitely additive. Let $\mathcal{C}(A)$ and $\mathcal{C}(B)$ be disjoint. Since every cylinder of rank $t_{0}$ is also a cylinder of rank $t$ for all $t\geq t_{0}$, we can choose $A,B$ to be subsets from the same $\mathcal{X}^{t}$. Thus, $A$ and $B$ are disjoint and $\mathcal{C}(A)\cup\mathcal{C}(B)=\mathcal{C}(A\cup B)$. So,
$P_{0}(\mathcal{C}(A\cup B))$ $=\sum_{x\in A\cup B}p(x)$
$=\sum_{x\in A}p(x)+\sum_{x\in B}p(x)$
$=P_{0}(\mathcal{C}(A))+P_{0}(\mathcal{C}(B)),$
@@
proving that $P_{0}$ is finitely additive, hence countably additive, hence a probability measure on $\mathscr{C}_{0}$. So, by the Extension Theorem [53], $P_{0}$ has a unique extension to a probability measure $P$ on $\mathscr{C}$. Define $\Phi_{23}(p)=P$.
@@
It is obvious that $(\Phi_{32}\circ\Phi_{23})(p)=p$ for all $p\in\mathbb{L}_{2}$. To prove that $(\Phi_{23}\circ\Phi_{32})(P)=P$ for all $P\in\mathbb{L}_{3}$, on the other hand, is quite subtle. Let $P_{1}\in\mathbb{L}_{3}$. Define $p=\Phi_{32}(P_{1})$ and $P_{2}=\Phi_{23}(p)$. While the uniqueness part of the Extension Theorem allowed us to define $\Phi_{23}$, it does not guarantee that $P_{1}=P_{2}$. To prove the latter, we will use Theorem 3.3 of [11], a consequence of Dynkin’s $\pi$-$\lambda$ theorem, which can be used to prove the uniqueness part of the Extension Theorem as well. Let $\mathscr{G}$ be the class of the empty set and thin cylinders. It is easy to check that $\mathscr{G}$ is closed under binary intersections and is, therefore, a $\pi$-system. Also, every cylinder can be decomposed into a finite union of thin cylinders. Thus, the algebra generated by $\mathscr{G}$ is $\mathscr{C}_{0}$, hence the $\sigma$-algebra generated by $\mathscr{G}$ is $\mathscr{C}$. So, by Theorem 3.3 of [11], any two probability measures on $\mathscr{C}$ that agree on $\mathscr{G}$ also agree on $\mathscr{C}$. Let $x\in\cup_{t=1}^{\infty}\mathcal{X}^{t}$. Then $P_{2}(\mathcal{C}(\{x\}))=p(x)=P_{1}(\mathcal{C}(\{x\}))$. So, $P_{1}$ and $P_{2}$ agree on $\mathscr{G}$, and so they are equal. Thus, $\Phi_{23}$ is the inverse of $\Phi_{32}$. ∎
@@
Note that the mapping $\Phi_{23}$ is standard. It follows the steps in [11] almost exactly until the definition of $P_{0}$ where [11] instead defines the product measure, which is the independent version of $P_{0}$. But, in the dependent setting, our $P_{0}$ is also standard, see e.g. Eq. (8) of [13], as it is the only possible definition that satisfies finite additivity since $\mathcal{C}(A)=\cup_{x\in A}\mathcal{C}(\{x\})$. The final step of extending $P_{0}$ to $P$ also follows as in [11]. But the subtle point that we have made in Proposition 6 is that for every $p\in\mathbb{L}_{2}$ there is a unique $P\in\mathbb{L}_{3}$ such that $P(\mathcal{C}(\{x\}))=p(x)$ for all $x\in\cup_{t=1}^{\infty}\mathcal{X}^{t}$. While the Extension Theorem guarantees that the extension from $\mathscr{C}_{0}$ to $\mathscr{C}$ is unique, which implies that $\Phi_{23}$ is well-defined, it does not prove that the extension from $\mathscr{G}$ to $\mathscr{C}$ is unique. So, while previous works showed that $\Phi_{23}$ is well-defined, they did not seem to notice that it is a bijection.
@@
Propositions 5 and 6 have important philosophical implications. Specifically, assuming that the next word in a sequence of words is always governed by a probability distribution is equivalent to assuming that infinite sequences of words are governed by a distribution. Moreover, the underlying functions in the two cases are in one-to-one correspondence with one another. But, it is conceptually easier to believe that there is an underlying probabilistic language space of the first kind because we naturally think of a sequence of words along with its extra-mathematical context, such as body language and flowers along the road. On the other hand, an underlying probabilistic language space of the third kind implies that language is a stochastic process. To see this, note that every probability measure is the pushforward measure of some random variable: just let $X$ be the identity mapping from $(\Omega,\mathscr{F})$ to itself, then $PX^{-1}=P$. In particular, for every $P\in\mathbb{L}_{3}$, there is a probability space $(\Omega,\mathscr{F},\mu)$ and a $(\mathcal{X}^{\mathbb{N}},\mathscr{C})$-valued random variable $X$ on $(\Omega,\mathscr{F},\mu)$ such that $P=\mu X^{-1}$. Each $X_{t}$ is a $(\mathcal{X},2^{\mathcal{X}})$-valued random variable, so $X$ is a stochastic process. Furthermore, its finite-dimensional distributions $\mu(X_{1},\ldots,X_{t})^{-1}$ are given by the probability mass function $\Phi_{32}(P)$ for each $t\in\mathbb{N}$.
For the rest of the paper, we will implicitly assume an underlying stochastic sequence while only using the lower-case $p$ for fitting a model to it. Thus, we can think of $p(x)$ as the probability that the stochastic sequence up to index $|x|$ is $x$, i.e. $p(x)=P(\mathcal{C}(\{x\}))$. And we can think of $p(y\mid x)$ as the probability that the value of the stochastic sequence at index $|x|+1$ is $y$ conditioned on the stochastic sequence up to index $|x|$ being $x$, i.e. $p(y\mid x)=P(\mathcal{C}(\{(x,y)\}))/P(\mathcal{C}(\{x\}))$. Here, we are making the usual abuse of notation where $p$ is used to denote both $p$ and $\Phi_{12}(p)$. We do this because the |||| makes it clear whether the function is in $\mathbb{L}_{1}$ or $\mathbb{L}_{2}$. To summarize, if we start with or define $p\in\mathbb{L}_{1}$, then we use $p$ to denote $\Phi_{12}(p)$ as well, and we can interpret the meaning of both in terms of $\mathbb{L}_{3}$.